# Perturbation-based Effect Measures for Compositional Data

### v2

@@ 1.2 Notation
We use superscripts to index coordinates of a vector throughout, that is, for $x\in\mathbb{R}^{d}$ we write $x^{j}$ for the $j$th coordinate of $x$ and denote by $e_{j}$ the $j$th canonical unit vector in $\mathbb{R}^{d}$. For all $n\in\mathbb{N}$, we define $[n]:=\{1,\dots,n\}$. For a real-valued function $f:\mathcal{X}\rightarrow\mathbb{R}$ with $\mathcal{X}\subseteq\mathbb{R}^{d}$, we denote by $\partial_{x^{j}}$ the partial derivative of $f$ with respect to the $x^{j}$ coordinate and the gradient (vector of partial derivatives) of $f$ is denoted by $\nabla f$. When differentiating a function of a single variable at a boundary point of the domain, we only consider the one-sided limit. For $p\geq 1$ and $x\in\mathbb{R}^{d}$, we define $\|x\|_{p}:=\left(\sum_{j=1}^{d}\left|x^{j}\right|^{p}\right)^{1/p}$. Lastly, we define $\mathsf{C}(z):=z/\sum_{j=1}^{d}z^{j}$, the (simplex) closure operator, with the additional convention that for $0\in\mathbb{R}^{d}$ we have $\mathsf{C}(0)=(1/d,\dots,1/d)=:z_{\mathrm{cen}}$.
@@ 2 Target parameters via perturbations
    $v_{\psi}(z):=\mathds{1}_{\{\omega_{\psi}(z)\neq 0\}}\frac{\omega_{\psi}(z)}{\|\omega_{\psi}(z)\|_{1}}\quad\text{}\quad s_{\psi}(z):=\|\omega_{\psi}(z)\|_{1},$
@@
Proposition 1 (derivative-isolating).
@@ 2.1 Perturbations on the simplex
We now turn to the construction of perturbations on the simplex. For binary perturbations, we only need to specify the endpoint of the perturbation, while for directional perturbations it suffices to choose a direction and a speed to construct a perturbation of the form (1). For a fixed endpoint there is a canonical (potentially geometry-dependent) choice of direction that is given by the straight line from a point to its endpoint. The choice of speed is, however, more subtle and it should be selected in a way that provides a clear interpretation. We propose two general approaches for selecting speeds: (a) by directly specifying an interpretable speed and (b) by starting from an interpretable summary statistic, which decreases when moving away from the endpoint, and interpreting the resulting effect as the effect on the response of slightly increasing the value of the statistic. For both approaches, Proposition 2 provides – under additional assumptions – explicit derivative-isolating reparametrizations, which can then be used for efficient estimation.
In the remainder of this section, we discuss interpretable simplex endpoints in Section 2.1.1, provide details on the selection of directions and speed together with the construction of derivative-isolating reparametrizations in Section 2.1.2 and finally in Section 2.1.3 discuss how one could construct directional perturbations that respect a different geometry than the Euclidean one. The perturbations introduced here are used in Section 2.2 to construct specific target parameters based on potential applications. The more applied reader may jump directly to that section and skip the more technical discussion on perturbations here.
@@ 2.1.1 Endpoints in the simplex
A second type of endpoint function, that can be non-constant, is given by mapping each $z\in\Delta^{d-1}$ to amalgamations on the simplex. An amalgamation is a well-studied operation (e.g., Aitchison, 1986) which consist of merging parts of the compositional components into others. Formally, we define the $(A\to B)$-amalgamation operator $\mathsf{A}_{A\to B}:\Delta^{d-1}\to\Delta^{d-1}$ for all $z\in\Delta^{d-1}$ by
@@
The $(A\to B)$-amalgamation has the effect of setting components in $A$ to 00, scaling up the components in $B$ such that the subcomposition in $B$ remains fixed (i.e., $\mathsf{C}(\mathsf{A}_{A\to B}(z)^{B})=\mathsf{C}(z^{B})$) and leaving the remaining components fixed. We call all perturbations $\psi$ with endpoint function $E_{\psi}:z\mapsto\mathsf{A}_{A\to B}(z)$ amalgamating perturbations. An important class of amalgamating perturbations are those where $A=[d]\setminus\{j\}$ and $B=\{j\}$ so that $\mathsf{A}_{A\to B}(z)=e_{j}$, i.e. perturbations that push towards a vertex of the simplex. Amalgamations, however, do not need to have fixed endpoints but allow for more general endpoints that can also depend on the initial point. Several perturbations on $\Delta^{2}$ are visualized in Figure 2 below.
@@ 2.1.2 Directions, speeds and derivative-isolating reparametrizations
Proposition 2.
@@ 2.1.3 Perturbations arising from alternative simplex geometries
Figure 2: Plots of (left) unit speed perturbations, (middle) alternative speed parametrizations (the diversity perturbation in blue has Gini-speed (see (10)) and multiplicative speed (see (8)) for the amalgamating perturbations in red and green) and (right) log-ratio perturbations for two different initial points $z_{1}$ and $z_{2}$. The types of perturbation are distinguished by color. The dashes represent the speed of the perturbation and are equally spaced in the $\gamma$-domain, so that longer lines correspond to faster speeds. The log-ratio diversity perturbation is not a straight line and does not have the same direction as the centering perturbations.
@@
Suppose we are interested in the effect of diversifying $z\in(\Delta^{d-1})^{\circ}$. As in Example 5, we can think of such effects as moving towards the center $z_{\mathrm{cen}}$, but now using the Aitchison geometry this corresponds to moving $\mathrm{CLR}(z)$ towards $\mathrm{CLR}(z_{\mathrm{cen}})=0$ along a straight line in CLR-space. This line transformed back to the simplex can be written as
$\{\mathrm{CLR}^{-1}((1-\gamma)\mathrm{CLR}(z))\mid\gamma\in[0,1]\}.$
Therefore, computing the diversifying direction in the Aitchison geometry amounts to differentiating the expression above at $\gamma=0$ and normalizing. Applying the chain rule, this yields
$\partial_{\gamma}\mathrm{CLR}^{-1}((1-\gamma)\mathrm{CLR}(z))^{j}\bigm{|}_{\gamma=0}=\sum_{k\in[d]:k\neq j}z^{j}z^{k}\log\left(\frac{z^{k}}{z^{j}}\right),$
which can be normalized to a unit $1$-norm vector. The direction of this derivative can be vastly different from the centering direction used in Example 5. To obtain a derivative-isolating reparametrization, we can instead define a unit-speed perturbation on the log-ratio scale (normalizing with the Euclidean norm as is natural here), $\widetilde{\psi}:(x,\gamma)\mapsto x-\gamma\frac{x}{\|x\|_{2}}$, and find a reparametrization on this scale. It can be shown that on the log-ratio scale, a derivative-isolating reparametrization is given by $\phi^{L}:x\mapsto-\|x\|_{2}$ and $\phi^{W}(z):=x/\|x\|_{2}$ and we can transform these to reparametrizations of $Z$ by composing with the $\mathrm{CLR}$-transform.
@@
Suppose, as in Example 3, we are interested in the effect of increasing the value of a particular component $z^{j}$ of $z\in(\Delta^{d-1})^{\circ}$. For $\mathrm{CLR}(z)\in\mathbb{R}^{d}$ we have a natural notion of increasing $z^{j}$ inherited from $\mathbb{R}^{d}$, namely adding $\gamma$ to $\mathrm{CLR}(z)^{j}$. The derivative of the resulting perturbation on the simplex is for all $k\in[d]$ given by
$\partial_{\gamma}\mathrm{CLR}^{-1}(\mathrm{CLR}(z)+\gamma e_{j})^{k}\bigm{|}_{\gamma=0}=\mathds{1}_{\{k=j\}}z^{j}-z^{j}z^{k}.$
@@ 2.2.1 Effects of changes in individual components
As a first example we consider perturbation effects that capture how a response is associated with changes in individual components. These types of effects are, for example, useful as a variable selection tool in compositional data. In Section 4.3, we illustrate this for a human gut microbiome study. For each study participant the body mass index (BMI) $Y$ and their gut microbiome composition $Z$ are measured. We then want to determine for which microbes a high abundance is associated with a high BMI. The simplex constraint implies that for all coordinates $j$ the conditional independence
$Y\mbox{${}\perp\mkern-11.0mu\perp{}$}Z^{j}\,|\,Z^{-j}$ (11)
is trivially satisfied and feature importance or influence measures based on (11) are meaningless. Intuitively, this stems from the fact that changing a single component always requires a change in at least some of the remaining components. We therefore propose – for a fixed component $j$ – to consider perturbations that increase the $j$th component while decreasing all remaining components proportionally. We define the compositional feature influence for the $j$th feature or simply $\mathrm{CFI}^{j}$ as $\tau_{\psi}$ for any directional perturbation with endpoint $e_{j}$. To that end, we already considered these perturbations in Example 3 where we proposed two possible speeds leading to slightly different interpretations of the average perturbation effects: $\mathrm{CFI}^{j}_{\mathrm{unit}}$ is the perturbation that moves at a constant (unit) speed towards the vertex and $\mathrm{CFI}^{j}_{\mathrm{mult}}$ is the perturbation that moves at what we call multiplicative speed (see Example 3 for more intuition on these speeds). $\mathrm{CFI}_{\mathrm{mult}}^{j}$ has previously been proposed by Huang et al. (2023), however without an efficient estimation procedure. Huang et al. (2023) also prove that when a log-contrast regression model is well-specified the $j$th coefficient in the model is equal to $\mathrm{CFI}_{\mathrm{mult}}^{j}$ (this proof is also included in Section S6 of the supplementary material). A crucial difference between unit speed and multiplicative speed perturbations is how they account for observations with $Z^{j}=0$. As the multiplicative speed is 00 in this case such points do not contribute to $\mathrm{CFI}^{j}_{\mathrm{mult}}$, however, they do contribute to $\mathrm{CFI}^{j}_{\mathrm{unit}}$.
@@ 3.1 Efficient estimation using semiparametric theory
We now give a brief introduction to semiparametric estimation of statistical functionals and, in particular, we describe how naive plug-in estimation of target parameters based on nonparametric methods rarely leads to efficient tests and how to remedy this. This perspective is not new and goes back to at least Pfanzagl (1982) with several more recent monographs and reviews (Bickel et al., 1993; Tsiatis, 2006; Chernozhukov et al., 2018; Kennedy, 2023). For illustration, we focus on the estimation of $\tau_{\psi}$, that is, the estimation of
@@
Figure 3: Plot of $95\%$ confidence intervals for a plug-in, one-step and cross-fit one-step estimator of $\lambda_{\psi}$ across 100 repetitions in the partially linear, binary simulation setting described in Section 4.1 with $d=15$ and $n=1000$. The black dashed line indicates the true value of the parameter and we highlight the confidence intervals that do not contain the true parameter. The confidence intervals are ordered first by whether the true value is in the interval, then by the value of the estimate.
@@ 4 Numerical experiments
We now present three numerical experiments illustrating the usage and performance of our methods. In Section 4.1 we investigate the performance of the semiparametric estimators discussed in Section 3. In Section 4.2.1 we analyze the effect of racial diversity on income on a semi-synthetic dataset based on data from the 1994 US census. In Section 4.2.2, we analyze the effect of racial diversity on grades in New York schools. In Section 4.3 we investigate variable influence (importance) measures for the prediction of BMI from gut microbiome data. The code for all of our experiments is available at https://github.com/ARLundborg/PerturbationCoDa. For all regressions we use the scikit-learn package (Pedregosa et al., 2011) except for the $\ell^{1}$-penalized log-contrast regression in Section 4.3 where we used the c-lasso package (Simpson et al., 2021).
4.1 Semiparametric estimators
  • NPM_no_x: The same estimators as NPM but without cross-fitting.
  • PLM_no_x: The same estimators as PLM but without cross-fitting.
The variance of the NPM and PLM estimates is computed based on the provided algorithms for the NPM and PLM methods while the variance for the plug-in is estimated using an estimator akin to the one given in (14). Each estimator above is itself based on multiple of the regression estimates $\widehat{f}$, $\widehat{\pi}$, $\widehat{\rho}$, $\widehat{g}$ and $\widehat{m}$. For the regression estimates, we use a $5$-fold cross-validation based estimator which for each fit selects among the following three models: a simple model that predicts the mean of the outcome (or the proportion of each class for binary classification), a random forest with $250$ trees with trees grown to a maximum depth of $2$ and a random forest with $250$ trees and no depth restriction. For the score estimator we use the location-scale based estimator described in Section S3.4 of the supplementary material, that uses the same $5$-fold cross-validated regression estimator for the location and scale nuisance regressions.
  • Binary $L$ with partially linear $Y$:
  • Binary $L$ with nonparametric $Y$:
  • Continuous $L$ with partially linear $Y$:
    $L=B(W_{1})+\xi\quad\text{and}\quad Y=L+B(W_{1})+\varepsilon$
  • Continuous $L$ with nonparametric $Y$:
The noise variables $U_{0}\sim\mathrm{Bern}(0.8)$, $U_{1}\sim\mathrm{Bern}(0.5)$, $\varepsilon\sim\mathcal{N}(0,1)$ and $\xi\sim\mathcal{N}(0,1)$ are all sampled independently (and independent of $W$). The true value of $\lambda$ in the binary $L$ settings and $\tau$ in the continuous $L$ settings is $1$. For each setting we simulate data and compute an estimate with a corresponding asymptotic $95\%$ confidence interval using each of the methods described above. We repeat the experiment $100$ times and compute coverage rates of the $95\%$ confidence intervals for $d=15$ in Figure 4 while coverage rates for $d=3$ and $d=75$ are reported in Section S4.2 of the supplementary material.
Figure 4: Estimated coverage rates of the $95\%$ confidence intervals for each different estimation procedure over $100$ repetitions for each of the settings described in Section 4.1 with $d=15$. PLM works well if $Y$ follows a partially linear model, but can fail if this assumption is not satisfied as in the nonparametric $Y$ setting. NPM works in both settings but can be less robust for small sample sizes. Further, it can be seen that the plug-in estimator and the estimators without cross-fitting do not lead to nominal coverage in some of these settings.

### v3

@@ 1.2 Notation
We use superscripts to index coordinates of a vector throughout, that is, for $x\in\mathbb{R}^{d}$ we write $x^{j}$ for the $j$th coordinate of $x$ and denote by $e_{j}$ the $j$th canonical unit vector in $\mathbb{R}^{d}$. For all $n\in\mathbb{N}$, we define $[n]:=\{1,\dots,n\}$. For a real-valued function $f:\mathcal{X}\rightarrow\mathbb{R}$ with $\mathcal{X}\subseteq\mathbb{R}^{d}$, we denote by $\partial_{x^{j}}$ the partial derivative of $f$ with respect to the $x^{j}$ coordinate and the gradient (vector of partial derivatives) of $f$ is denoted by $\nabla f$. When differentiating a function of a single variable at a boundary point of the domain, we only consider the one-sided limit. For $p\geq 1$ and $x\in\mathbb{R}^{d}$, we define $\|x\|_{p}:=\left(\sum_{j=1}^{d}\left|x^{j}\right|^{p}\right)^{1/p}$. Lastly, we define $\mathsf{C}(z):=z/\sum_{j=1}^{d}z^{j}$, the (simplex) closure operator.
@@ 2 Target parameters via perturbations
    $v_{\psi}(z):=\mathds{1}_{\{\omega_{\psi}(z)\neq 0\}}\frac{\omega_{\psi}(z)}{\|\omega_{\psi}(z)\|_{1}}\quad\text{and}\quad s_{\psi}(z):=\|\omega_{\psi}(z)\|_{1},$
@@
Proposition 1 (Isolating derivatives).
@@ 2.1 Perturbations on the simplex
We now turn to the construction of perturbations on the simplex. For binary perturbations, we only need to specify the endpoint of the perturbation, while for directional perturbations it suffices to choose a direction and a speed to construct a perturbation of the form (1). For a fixed endpoint there is a canonical (geometry-dependent) choice of direction that is given by the straight line from a point to its endpoint. The choice of speed is, however, more subtle and it should be selected in a way that provides a clear interpretation. We propose two general approaches for selecting speeds: (a) by directly specifying an interpretable speed and (b) by starting from an interpretable summary statistic, which decreases when moving away from the endpoint, and interpreting the resulting effect as the effect on the response of slightly increasing the value of the statistic. For both approaches, Proposition 2 provides – under additional assumptions – explicit derivative-isolating reparametrizations, which can then be used for efficient estimation.
In the remainder of this section, we discuss interpretable simplex endpoints in Section 2.1.1, provide details on the selection of directions and speed together with the construction of derivative-isolating reparametrizations in Section 2.1.2 and finally in Section 2.1.3 discuss how one could construct directional perturbations that respect a different geometry than the Euclidean one. The perturbations introduced here are used in Section 2.2 to construct specific target parameters based on potential applications. Section 2.2 may be particularly interesting for the reader who is more interested in applications of our framework.
@@ 2.1.1 Endpoints in the simplex
A second type of endpoint function, that can be non-constant, is given by mapping each $z\in\Delta^{d-1}$ to amalgamations on the simplex. An amalgamation is a well-studied operation (e.g., Aitchison, 1986) which consists of adding together the values of groups of components of $z$ resulting in a lower dimensional composition. Formally, for disjoint $A,B\subseteq[d]$, we define the $(A\to B)$-amalgamation operator $\mathsf{A}_{A\to B}:\Delta^{d-1}\to\Delta^{d-1}$ for all $z\in\Delta^{d-1}$ by
@@
The $(A\to B)$-amalgamation has the effect of setting components in $A$ to 00, scaling up the components in $B$ such that the subcomposition in $B$ remains fixed (i.e., $\mathsf{C}(\mathsf{A}_{A\to B}(z)^{B})=\mathsf{C}(z^{B})$) and leaving the remaining components fixed. We call all perturbations $\psi$ with endpoint function $E_{\psi}:z\mapsto\mathsf{A}_{A\to B}(z)$ amalgamating perturbations. We give an example of a scientific question involving a general amalgamation in Section 2.2.3. An important class of amalgamating perturbations are those where $A=[d]\setminus\{j\}$ and $B=\{j\}$ so that $\mathsf{A}_{A\to B}(z)=e_{j}$, i.e. perturbations that push towards a vertex of the simplex. Amalgamations, however, do not need to have fixed endpoints but allow for more general endpoints that can also depend on the initial point. Several perturbations on $\Delta^{2}$ are visualized in Figure 2 below.
@@ 2.1.2 Directions, speeds and derivative-isolating reparametrizations
Proposition 2 (Finding derivative-isolating reparametrizations).
@@ 2.1.3 Perturbations arising from alternative simplex geometries
Figure 2: Plots of (left) unit speed perturbations, (middle) alternative speed parametrizations and (right) log-ratio perturbations for two different initial points $z_{1}$ and $z_{2}$. The types of perturbation are distinguished by color. The alternative speed for the diversity perturbation in blue is Gini-speed (see (10)) and for the amalgamating perturbations in red and green the speed is multiplicative speed (see (8)). The dashes represent the speed of the perturbation and are equally spaced in the $\gamma$-domain, so that longer lines correspond to faster speeds. The log-ratio diversity perturbation is not a straight line and does not have the same direction as the centering perturbations.
@@
Suppose we are interested in the effect of diversifying $z\in(\Delta^{d-1})^{\circ}$. As in Example 5, we can think of such effects as moving towards the center $z_{\mathrm{cen}}$, but now using the Aitchison geometry this corresponds to moving $\mathrm{CLR}(z)$ towards $\mathrm{CLR}(z_{\mathrm{cen}})=0$ along a straight line in CLR-space. Using the Aitchison geometry, this defines a perturbation via
$\psi_{\mathrm{Ait}}:(z,\gamma)\mapsto\mathrm{CLR}^{-1}((1-\gamma)\mathrm{CLR}(z)).$
We can compute $\partial_{\gamma}\psi_{\mathrm{Ait}}(z,\gamma)\bigm{|}_{\gamma=0}$ to find a perturbation $\psi$ of the form (1) for which $\tau_{\psi}=\tau_{\psi_{\mathrm{Ait}}}$. Applying the chain rule, we obtain
$\partial_{\gamma}\psi_{\mathrm{Ait}}(z,\gamma)^{j}\bigm{|}_{\gamma=0}=\sum_{k\in[d]:k\neq j}z^{j}z^{k}\log\left(\frac{z^{k}}{z^{j}}\right),$
which can be decomposed into a speed and direction as in Definition 1 to obtain $\psi$. The direction of this derivative can be vastly different from the centering direction used in Example 5. To obtain a derivative-isolating reparametrization, since $\psi_{\mathrm{Ait}}$ shrinks towards 00 in CLR-space, it is straightforward to show that $\phi^{L}:z\mapsto-\|\mathrm{CLR}(z)\|_{2}$ and $\phi^{W}:z\mapsto\mathrm{CLR}(z)/\|\mathrm{CLR}(z)\|_{2}$ are derivative-isolating reparametrizations.
@@
Suppose, as in Example 3, we are interested in the effect of increasing the value of a particular component $z^{j}$ of $z\in(\Delta^{d-1})^{\circ}$. There is a natural notion of increasing $z^{j}$ inherited from CLR-space, namely adding $\gamma$ to $\mathrm{CLR}(z)^{j}$ and transforming back to the simplex. The resulting perturbation is given by
$\psi:(z,\gamma)\mapsto\mathrm{CLR}^{-1}(\mathrm{CLR}(z)+\gamma e_{j}).$
The derivative of the resulting perturbation is for all $k\in[d]$ given by
$\partial_{\gamma}\psi(z,\gamma)^{k}\bigm{|}_{\gamma=0}=\mathds{1}_{\{k=j\}}z^{j}-z^{j}z^{k}.$
@@ 2.2.1 Effects of changes in individual components
As a first example we consider perturbation effects that capture how a response is associated with changes in individual components. These types of effects are, for example, useful as a variable selection tool in compositional data. In Section 4.2, we illustrate this for a human gut microbiome study. For each study participant the body mass index (BMI) $Y$ and their gut microbiome composition $Z$ are measured. We then want to determine for which microbes a high abundance is associated with a high BMI.The simplex constraint implies that for all coordinates $j$ we have that $Y$ is independent of $Z^{j}$ given $Z^{-j}$ hence any feature importance or influence measure based on measuring this conditional dependence is meaningless.
Intuitively, this stems from the fact that knowing $Z^{-j}$ implies that we also know $Z^{j}$ via the relation $Z^{j}=1-\|Z^{-j}\|_{1}$ hence $Z^{j}$ can never reveal any additional information regarding $Y$. We therefore propose – for a fixed component $j$ – to consider perturbations that increase the $j$th component while decreasing all remaining components proportionally. We define the compositional feature influence for the $j$th feature or simply $\mathrm{CFI}^{j}$ as $\tau_{\psi}$ for any directional perturbation with endpoint $e_{j}$. To that end, we already considered these perturbations in Example 3 where we proposed two possible speeds leading to slightly different interpretations of the average perturbation effects: $\mathrm{CFI}^{j}_{\mathrm{unit}}$ is the perturbation that moves at a constant (unit) speed towards the vertex and $\mathrm{CFI}^{j}_{\mathrm{mult}}$ is the perturbation that moves at what we call multiplicative speed (see Example 3 for more intuition on these speeds). $\mathrm{CFI}_{\mathrm{mult}}^{j}$ has previously been proposed by Huang et al. (2023), however without an efficient estimation procedure. Huang et al. (2023) also prove that when a log-contrast regression model is well-specified the $j$th coefficient in the model is equal to $\mathrm{CFI}_{\mathrm{mult}}^{j}$ (this proof is also included in Section S6 of the supplementary material). A crucial difference between unit speed and multiplicative speed perturbations is how they account for observations with $Z^{j}=0$. As the multiplicative speed is 00 in this case such points do not contribute to $\mathrm{CFI}^{j}_{\mathrm{mult}}$, however, they do contribute to $\mathrm{CFI}^{j}_{\mathrm{unit}}$.
@@ 3.1 Efficient estimation using semiparametric theory
We now give a brief introduction to semiparametric estimation of statistical functionals and, in particular, we describe how naive plug-in estimation of target parameters based on nonparametric methods rarely leads to efficient tests and how to remedy this. This perspective is not new and goes back to at least Pfanzagl (1982) with several more recent monographs and reviews (Bickel et al., 1993; Tsiatis, 2006; Chernozhukov et al., 2018; Kennedy, 2023). The recent surge in the popularity of these methods is fueled by the advent of powerful machine learning methods that use flexible black box models to learn the nuisance functions involved in the estimation of functionals of interest. For illustration, we focus on the estimation of $\tau_{\psi}$, that is, the estimation of
@@
Figure 3: Plot of $95\%$ confidence intervals for a plug-in, one-step and cross-fit one-step estimator of $\lambda_{\psi}$ across 100 repetitions in the partially linear, binary simulation setting described in Section S4.2 of the supplementary material with $d=15$ and $n=1000$. The black dashed line indicates the true value of the parameter and we highlight the confidence intervals that do not contain the true parameter. The confidence intervals are ordered first by whether the true value is in the interval, then by the value of the estimate.
@@ 4 Numerical experiments
We now present numerical experiments illustrating the usage and performance of our methods. In Section 4.1.1 we analyze the effect of racial diversity on income on a semi-synthetic dataset based on data from the 1994 US census. In Section 4.1.2, we analyze the effect of racial diversity on grades in New York schools. In Section 4.2 we investigate variable influence (importance) measures for the prediction of BMI from gut microbiome data. In addition to the experiments included here, Section S4.2 of the supplementary material contains an investigation of the performance of the semiparametric estimators discussed in Section 3 when the conditioning variable $W$ is compositional. The code for all of our experiments is available at https://github.com/ARLundborg/PerturbationCoDa. For all regressions we use the scikit-learn package (Pedregosa et al., 2011) except for the $\ell^{1}$-penalized log-contrast regression in Section 4.2 where we used the c-lasso package (Simpson et al., 2021).